# Multimodal MRI-based Detection of Amyloid Status in Alzheimer’s Disease Continuum

### v1

@@ 4.3 Feature relevance & Statistical analysis
Figure 3A shows violinplots for the top sMRI features, while Table 3a shows the corresponding percentages of mean GBP contribution across A positive correctly classified patients for the same regions, weighted by their volumes. The Table also reports the p-values and FDR-corrected p-values of the statistical tests performed on the input features. We identified significant differences between classes after FDR correction in the precuneus, cingulate gyrus (posterior division), thalamus, hippocampus, supracalcarine and intracalcarine cortices, amygdala, and temporal occipital fusiform cortex showing an overall comparison direction of .
Figure 3B shows violinplots representing the graph node strength for the top rs-fMRI nodes, while Table 3b shows the mean percentages of GBP contribution extracted from the same nodes (ICs) across A positive patients. After FDR correction, the precuneus (IC48) was the only region with significant differences (). Before correction, another precuneus component (IC43) and precentral gyrus (IC14) were also significant with , while the left inferior parietal lobule (IC38) exhibited a trend towards uncorrected statistical significance (p, with ).
Figure 3C shows violinplots representing the graph betweenness centrality for the top dMRI nodes, and Table 3c shows the mean percentages of GBP contribution for the same nodes across A positive correctly classified patients. The precuneus LH and cerebellum RH were significant when considering uncorrected p-values (), but none of the results survived multiple comparison corrections. Moreover, the right superior parietal gyrus showed a trend towards significance (p, ).
@@ 5 Discussion
In this work, we presented a multimodal data fusion framework that integrates multiple MRI techniques (sMRI, rs-fMRI, and dMRI) for the classification of A status. A post-hoc XAI analysis and evaluation was performed using GBP to identify feature importance, complemented by statistical analyses on the identified input features for validation via post-hoc plausibility assessment. The underlying assumption is that maximally contributing regions would bring discriminative information across the target classes.
@@ 5.1 Classification performance
Our results showed that the multimodal framework outperformed single-modality models in terms of classification performance, in particular for rs-fMRI and dMRI, while the sMRI achieved performance close to the multimodal one, but with high variance across the evaluation folds. This confirms the added value provided by multimodality approaches which improve classification accuracy and, at the same time, carry more information, particularly relevant for the study of complex neurodegenerative diseases influenced by multi-domain factors.
@@
Table 4 shows the performance of state-of-the-art works for the classification of A positive and negative subjects. Of note, current state-of-the-art works use multiple datasets and approaches for addressing this task without considering a common benchmark dataset. Hence, the comparison of the different approaches is imperfect and any resulting conclusions held tentatively. Most of the works relied on PET scans to address this task, e.g., [14] developed a CNN model that takes in input the different views (axial, coronal, and sagittal) of a 3D volume; while [6] used three different well-known 3D-CNN (Inception3D, ResNet3D, and VGG3D) to address this task using the full 3D volume of PET scans as input. They achieved an accuracy of around on average in the test set, with a maximum accuracy of by [6] using the VGG3D architecture. In the same way, [37] employed a 3D-CNN to identify the A status reporting both the average validation accuracy and also the accuracy on a hold-out test set (ADNI data) using the ensemble method created with the best models across the different folds, achieving accuracies of around . Recently, [16] employed different combinations of feature selection methods and classifiers in order to analyze FDG-PET-derived features. They achieved an AUC of employing LASSO with the GNB model. On the other hand, only few works have used MRI modalities to classify A positive and negative subjects. [18] used only the hippocampus region for predicting A positivity, testing their network on two different classification tasks: AD A POS versus CN A NEG, and MCI A POS versus MCI A NEG. They achieved an accuracy of and , in the first and second tasks, respectively. Lastly, [17] implemented a 3D-CNN for detecting A status from sMRI images defining the A status of the CN, MCI, and AD patients considering the PET cortical standardized uptake value ratio. They achieved an accuracy of .
@@
Comparing our proposed method with these five works, we outperformed the method that relied on 3D FDG-PET proposed by [14] and the performance of [17]. Our study achieved competitive performance in the validation set compared to [6] using amyloid PET images, although our framework achieved a higher accuracy compared to the ResNet3D they used, and a similar accuracy to the one reached by the model proposed by [18] using the hippocampus region only. While the current classification of A positive or negative subjects can only be performed through CSF tests or amyloid PET scans with high precision, the development of different approaches based on non-invasive MRI modalities for determining A status could improve patient care highlighting specific mechanisms that underlie AD. We further showed the ability of our model to successfully integrate structural GM features extracted from a 3D-CNN along with functional and structural connectivity analyzed by two types of GCNs. Notably, while a few studies have recently started to explore SC/FC in conjunction with GCNs in the AD classification tasks ([38, 39]), results are still limited in this specific problem calling for further investigation in the AD continuum.
@@ 5.2 Explainability analysis
The post-hoc analysis, performed on the correctly classified subjects, consisted of two steps: i) extraction of GBP attribution maps for A positive patients and corresponding percentage of contribution for each brain region both within and between modalities for the average positive patient; ii) statistical analysis on the input features of the most important regions/nodes identified by GBP.
@@
In our analysis of sMRI attribution maps, multiple brain regions involved in AD neurodegeneration were identified. These regions had a higher percentage of contribution relative to the others. These regions belonged to subcortical areas (hippocampus, thalamus, putamen, accumbens, and amygdala) and temporal/occipital/parietal areas (posterior cingulate gyrus, precuneus cortex, supracalcarine and intracalcarine cortices, and the temporal occipital fusiform cortex). The hippocampus, in particular, is a well-known biomarker for AD that is subject to high levels of atrophy. Some studies have suggested that this atrophy is attributable to the deposition of A plaques ([40]). An important region that is connected with the hippocampus is the cingulate gyrus which showed a strong reduction of GM in AD patients ([41, 42]). In a clinical study performed by [43], significant associations were detected between A accumulation and GM atrophy in the hippocampus and posterior cingulate gyrus for the MCI and AD A+ patients. Moreover, other subcortical structures, like the thalamus and putamen, have a high atrophy rate in clinical patients relative to healthy individuals ([44, 43]), probably due to A deposition. A previous study highlighted an increase in SUVR (derived from florbetapir PET amyloid) in both putamen and thalamus in the preclinical stages of AD ([45]). In the statistical analysis performed on the input GM volumes, significant differences were detected in the hippocampus, posterior cingulate gyrus, amygdala, thalamus, and precuneus, which demonstrated a considerable decrease of GM in the A positive patients. These findings were consistent with preexisting literature. Additionally, [46] highlighted other regions (that we also found to be relevant) involved in the different phases of A deposition (i.e., proisocortex, allocortical areas, diencephalic nuclei, and striatum). For the rs-fMRI, the brain regions (ICs) that most contributed to the final classification resided in the DM and CC networks, along with three regions for VI and SM networks (two and one, respectively). The DM network is involved in memory, self-knowledge, and thinking, and it is highly related to AD ([8, 41]) Previous studies also suggested that DMN is associated with A plaques ([47]). The CC network is associated with selective attention, working memory, and stimulus-response mapping ([48, 49]). The statistical analysis, performed on the node strength values, detected one region with significant differences between the two groups after FDR correction and three regions before (uncorrected p-values). One more region, the left inferior parietal lobule, exhibited a trend toward significance. Specifically, the A positive class had a significantly increased node strength in the precentral gyrus (IC14, SM network), which was consistent with the findings of [50] and [51]. The precuneus regions (IC43 and IC48, DM network) were also significantly different between groups with increased strength in the positive group. As in other works ([52, 53]), this suggested that the precuneus plays a key role and undergoes pathological changes related to AD. Previous works also identified the posterior DM (the precuneus and posterior cingulate cortex) ([9]) and parts of the CC network (i.e., the middle frontal gyrus and hippocampus) ([54, 55]) as strongly affected by A deposition. Lastly, in the dMRI channel, GBP identified multiple cortical and subcortical regions as relevant to the final predictions of our model. The cerebellum was also revealed to be important by the model. Of note, the rostral middle frontal gyrus was assigned a high percentage of contribution relative to the other regions. Additionally, the middle temporal gyrus, precuneus, thalamus, and superior parietal regions showed high importance relative to the other regions. After FDR correction, no region had statistically significant group differences in betweenness centrality values. On the other hand, based upon uncorrected p-values, significant differences (p) were highlighted in the precuneus and cerebellum (right hemisphere), while the superior parietal region was close to significance (p=). Regarding the two nodes with significant uncorrected differences, the precuneus, which is also identified in the rs-fMRI, is an important hub for functional operations because it is highly connected with other regions by both short- and long-range WM fibers ([53]). On the other hand, different studies highlighted how the cerebellum is subject to increase A deposition in AD pathology ([56, 57]) at different stages (stage in [58] and stage in [46]) relative to normally aged subjects. Widespread structural alterations in WM tracts connecting cortical and subcortical regions in both hemispheres were detected between A negative CN and positive preclinical AD patients by [59]. Finally, when combining the explanations of all channels, we found that sMRI most contributed to the final prediction. This was expected as we initially assumed that the effects of A plaques would predominantly impact GM volumes.
@@ 5.3 Common regions across modalities
Multiple brain regions were identified as important by GBP across all three channels. Specifically, the precuneus was highlighted in all three modalities. It was one of the most interesting regions as it plays a pivotal role in the transmission of functional information due to the high concentration of WM tracts linking the precuneus with other brain areas. Our finding of decreased patient GM and increased functional node strength for the region could be related to the functional compensation effects of AD and merits further investigation ([52, 53]). Additionally, the cingulate gyrus (posterior division), calcarine gyrus, and hippocampus, were important in both sMRI and rs-fMRI. Only the thalamus was relevant to both sMRI and dMRI, and the middle frontal gyrus was the only area important to both rs-fMRI and dMRI. Interestingly, the middle frontal gyrus was a region highlighted in both rs-fMRI and dMRI, but not in the sMRI. This further underlines the importance of integrating different (and complementary) modalities to provide a more complete picture of the complex pathological mechanisms.
@@ 5.4 Main contributions and outcomes
We proposed a multimodal and explainable neuroimaging DL model for the classification of amyloid- positive or negative status. The main contribution of this work is a framework that i) successfully integrates volumetric features and connectivity information from sMRI, rs-fMRI, and dMRI data, extracted from one 3D-CNN and two GCNs, respectively, ii) obtains competitive classification performance (ACC=) relative to the state-of-the-art with the advantage of avoiding amyloid PET, iii) provides insight into the brain regions that most contribute to final model predictions using a GBP-based post-hoc analysis. Finally, we evaluated the XAI outcomes by assessing the discriminative power of the selected input features across classes.
With these analyses, we identified multiple brain regions that suffered from the effects of A plaque deposition, resulting in atrophy, functional changes, and WM changes relative to the control class. Furthermore, the analysis identified common regions across modalities, including the precuneus, hippocampus, thalamus, cingulate gyrus, calcarine gyrus, and middle frontal gyrus, strengthing the evidence of their involvement in this pathological process.
Our findings highlight the utility of MRI for studying the effects of A deposition and the importance of integrating complementary information to enable a better understanding of the differences related to amyloid status. However, we acknowledge that a main limitation of this work is the limited number of subjects in the dataset having both A information and all the considered MRI modalities. Further studies are thus needed to evaluate the generalizability of this multimodal framework to larger datasets, also including subjects with opposite A status (e.g., CN pos).

### v2

@@ 4.3 Feature relevance & Statistical analysis
Figure 3A shows the violinplots for the top sMRI features, while Table 3a shows the corresponding percentages of mean GBP contributions across A positive correctly classified subjects for the same regions, weighted by their volumes. The Table also reports the p-values and FDR-corrected p-values of the statistical tests performed on the input features. We identified significant differences between classes after FDR correction in the precuneus, cingulate gyrus (posterior division), thalamus, hippocampus, supracalcarine and intracalcarine cortices, amygdala, and temporal occipital fusiform cortex showing an overall comparison direction of .
Figure 3B shows violinplots representing the graph node strength for the top rs-fMRI nodes, while Table 3b shows the mean percentages of GBP contributions extracted from the same nodes (ICs) across A positive patients. After FDR correction, the precuneus (IC48) was the only region with significant differences (). Before correction, precuneus component (IC43) and precentral gyrus (IC14) were also significant with , while the left inferior parietal lobule (IC38) exhibited a trend towards uncorrected statistical significance (p, with ).
Figure 3C shows the violinplots representing the graph betweenness centrality for the top dMRI nodes, and Table 3c shows the mean percentages of GBP contributions for the same nodes across A positive correctly classified subject. The precuneus LH and cerebellum RH were significant when considering uncorrected p-values (), but none of the results survived multiple comparisons correction. Moreover, the right superior parietal gyrus showed a trend towards significance (p, ).
@@ 5 Discussion
In this work, we proposed a multimodal data fusion framework that integrates multiple MRI techniques (sMRI, rs-fMRI, and dMRI) for the classification of A status in an unbalanced cohort of subjects. A post-hoc XAI analysis and evaluation was performed using GBP to assess feature importance, complemented by statistical analyses on the so-identified input features for validation via post-hoc plausibility assessment.
@@ 5.1 Classification performance
Results showed that the multimodal framework outperformed single-modality models in terms of classification performance, in particular for rs-fMRI and dMRI, while the sMRI achieved performance close to the multimodal one, but with higher variance across the evaluation folds. This provides evidence of the added value brought by multimodality approaches in terms of classification accuracy while injecting complementary information shading light on the underlying neurophysiological mechanisms, which is particularly relevant for the study of complex neurodegenerative diseases influenced by multi-domain factors.
Table 4 shows the performance of state-of-the-art (SOA) works for the classification of A positive versus negative subjects. Of note, current SOA works use multiple datasets and approaches for addressing this task. It goes without saying that the lack of a common reference dataset inherently limits the relevance of performance comparison. However, reaching good accuracy and outperforming the SOA does not exhaust the contribution of the proposed approach, whose potential lies in decoding the neurophysiological changes induced by the A status as captured by the considered MRI modalities.
@@
Most of the SOA works relied on PET scans to address this task, e.g., [14] developed a CNN model that takes the different views (axial, coronal, and sagittal) of a 3D volume as input, while [6] used three different well-known 3D-CNN (Inception3D, ResNet3D, and VGG3D) models to address this task using the full 3D volume of PET scans. They achieved an accuracy of around on average in the test set, with a maximum accuracy of by [6] using the VGG3D architecture. In the same way, [39] employed a 3D-CNN to identify the A status reporting both the average validation accuracy and also the accuracy on a hold-out test set (ADNI data) using the ensemble method created with the best models across the different folds, achieving accuracies of around . Recently, [16] employed different combinations of feature selection methods and classifiers in order to analyze FDG-PET-derived features. They achieved an AUC of employing LASSO with the GNB model. Only a few works used MRI modalities to classify A positive and negative subjects. [18] focused on the hippocampus region only for detecting A positivity, testing the network on two different classification tasks: AD A positive versus CN A negative, and MCI A positive versus MCI A negative, achieving an accuracy of and , in the first and second tasks, respectively. Lastly, [17] implemented a 3D-CNN for detecting A status from sMRI images defining the A status of the CN, MCI, and AD patients considering the PET cortical standardized uptake value ratio reaching an accuracy of .
@@
Comparing the proposed method with these six works, we outperformed the method that relied on 3D FDG-PET proposed by [14] and the performance of [17]. Our study achieved competitive performance in the validation set compared to [6] using amyloid PET images, although our pipeline led to higher accuracy compared to the ResNet3D they used, and a similar accuracy to that in [18] using the hippocampus region only. Besides providing competitive performance, our model allows to successfully integrate structural GM features extracted from a 3D-CNN along with functional and structural connectivity relying on two ad-hoc GCNs. Notably, while a few studies have recently started to explore SC/FC in conjunction with GCNs in the AD classification task ([40, 41]), results are still limited, calling for further investigation in the AD continuum.
@@ 5.2 Explainability analysis
The post-hoc analysis, performed on the correctly classified subjects, consisted of two steps: i) extraction of GBP attribution maps for A mean positive subjects as well as the corresponding percentage of contribution for each brain region, both within and between modalities; ii) statistical analysis on the input features of the most important regions/nodes identified by GBP.
@@
The analysis of the sMRI attribution maps revealed that multiple brain regions involved in AD neurodegeneration could be identified. These regions had a higher percentage of contribution relative to the others and belonged to subcortical areas (hippocampus, thalamus, putamen, accumbens, and amygdala) and temporal/occipital/parietal areas (posterior cingulate gyrus, precuneus cortex, supracalcarine and intracalcarine cortices, and the temporal occipital fusiform cortex). The hippocampus, in particular, is a well-known biomarker for AD that is subject to high levels of atrophy. Some studies have suggested that this atrophy is attributable to the deposition of A plaques ([42]). An important region that is connected with the hippocampus is the cingulate gyrus which showed a strong reduction of GM in AD patients ([43, 44]). In a clinical study performed by [45], significant associations were detected between A accumulation and GM atrophy in the hippocampus and posterior cingulate gyrus for the MCI and AD A positive subjects. Some works also linked the nucleus accumbens region to AD pathology and progression. Evidence of GM loss and alteration of nucleus accumbens in MCI and AD, with respect to the CN, had previously been highlighted ([46, 47]). Additionally, [48] showed how A oligomers in nucleus accumbens can promote synaptic loss and motivation deficits in AD. Other subcortical structures, like the thalamus and putamen, have a high atrophy rate in clinical patients relative to healthy individuals ([49, 45]), probably due to A deposition. A previous study highlighted an increase in the standardized uptake value ratio (derived from florbetapir PET) in both putamen and thalamus in the preclinical stages of AD ([50]). Following the statistical analysis performed on the input GM volumes, significant differences were detected in the hippocampus, posterior cingulate gyrus, amygdala, thalamus, and precuneus, highlighting a considerable decrease of GM in the A positive subjects. These findings were consistent with preexisting literature. Additionally, [51] highlighted other regions (that we also found to be relevant) involved in the different phases of A deposition (i.e., proisocortex, allocortical areas, diencephalic nuclei, and striatum).
For the rs-fMRI, the brain regions (ICs) that most contributed to the final classification resided in the DM and CC networks, along with three regions for VI and SM networks (two and one, respectively). The DM network is involved in memory, self-knowledge, and thinking, and it is highly related to AD ([8, 43]). Previous studies also suggested that DMN is associated with the presence of A plaques ([52]). The CC network is associated with selective attention, working memory, and stimulus-response mapping ([53, 54]). The statistical analysis, performed on the node strength values, detected one region with a significant difference between the two groups after FDR correction and three regions before (uncorrected p-values). One more region, the left inferior parietal lobule, exhibited a trend toward significance. Specifically, the A positive class had a significantly increased node strength in the precentral gyrus (IC14, SM network), which was consistent with the findings of [55] and [56]. The precuneus regions (IC43 and IC48, DM network) were also significantly different between groups with increased strength in the positive group. As in other works ([57, 58]), this suggested that the precuneus plays a key role and undergoes pathological changes related to AD. Previous works also identified the posterior DM (the precuneus and posterior cingulate cortex) ([9]) and parts of the CC network (i.e., the middle frontal gyrus and hippocampus) ([59, 60]) as strongly affected by A deposition.
Lastly, in the dMRI channel, GBP identified multiple cortical and subcortical regions as relevant to the outcome of our model. The cerebellum resulted to be important for the model. Of note, the rostral middle frontal gyrus was assigned a high percentage of contribution relative to the other regions. Additionally, the middle temporal gyrus, precuneus, thalamus, and superior parietal regions showed high importance relative to the other regions. After FDR correction, no region had statistically significant group differences in betweenness centrality values. On the other hand, based upon uncorrected p-values, significant differences (p) were detected in the precuneus and cerebellum (right hemisphere), while the superior parietal region was close to significance (p=). Regarding the two nodes with significant uncorrected differences, the precuneus, which is also identified in the rs-fMRI, is an important hub for functional operations because it is highly connected with other regions by both short- and long-range WM fibers ([58]). On the other hand, different studies highlighted how the cerebellum is subject to increase A deposition in AD pathology ([61, 62]) at different stages (stage in [63] and stage in [51]) relative to normally aged subjects. Widespread structural alterations in WM tracts connecting cortical and subcortical regions in both hemispheres were detected between A negative CN and positive preclinical AD patients by [64].
When combining the explanations of all channels, we found that sMRI most contributed to the final prediction. This was expected as we initially assumed that the effects of A plaques would predominantly impact GM volumes.
@@ 5.3 Common regions across modalities
Multiple brain regions were identified as important by GBP across all three channels. Specifically, the precuneus was highlighted in all three modalities. It was one of the most interesting regions as it plays a pivotal role in the transmission of functional information due to the high concentration of WM tracts linking the precuneus with other brain areas. Our finding of decreased GM and increased functional node strength for the region in A positive subjects could be related to the functional compensation effects of AD and deserves further investigation ([57, 58]). Additionally, the cingulate gyrus (posterior division), calcarine gyrus, and hippocampus were important in both sMRI and rs-fMRI. Only the thalamus was relevant to both sMRI and dMRI, and the middle frontal gyrus was the only area important to both rs-fMRI and dMRI. Interestingly, the middle frontal gyrus was a region highlighted in both rs-fMRI and dMRI, but not in the sMRI. This further underlines the importance of integrating different (and complementary) modalities to provide a complete picture of the complex pathological mechanisms.
@@ 5.4 Main contributions and outcomes
We proposed a multimodal and explainable neuroimaging DL model for the classification of amyloid- positive or negative status. The main contribution of this work is a framework that i) successfully integrates volumetric features and connectivity information from sMRI, rs-fMRI, and dMRI data, extracted from one 3D-CNN and two GCNs, respectively, ii) obtains competitive classification performance (ACC=) relative to the state-of-the-art, iii) provides insight into the brain regions that most contribute to final model outcome using a GBP-based post-hoc analysis. Finally, we evaluated the XAI outcomes by assessing the discriminative power of the selected input features across classes.
With these analyses, we identified multiple brain regions that could be altered by A plaque deposition, resulting in atrophy, functional changes, and WM connectivity changes relative to the negative class. Furthermore, the analysis identified common regions across modalities, including the precuneus, hippocampus, thalamus, cingulate gyrus, calcarine gyrus, and middle frontal gyrus, strengthing the evidence of their involvement in this pathological process.
Our findings highlight the utility of MRI for studying the possible effects of A deposition and the importance of integrating complementary information to enable a better understanding of the differences related to amyloid status. However, we acknowledge that the two main limitations of this work are the limited number of subjects in the dataset having both A information and all the considered MRI modalities, and the unbalanced cohort including CN, SMC, and EMCI in the A negative group and EMCI, LMCI, and AD in the A positive one. Hence, the outcomes could be guided by those factors that were influenced by the amyloid accumulation as captured by the specific modality. In consequence, the GM branch captured atrophy, while the structural and functional connectivity branches captured the respective amyloid-induced connectivity modulations, leading to a combined effect due to A and dementia-related processes. Further studies are thus needed to assess the generalizability of the proposed multimodal framework to larger and possibly employing balanced datasets, relying on study cohorts composed of all clinical classes (CN, SMC, EMCI, LMCI, and AD) in both groups analyzed, A negative and positive individuals, in order to disentangle the contributions of the different factors and hence having the outcomes only related to A.